OpenStudy (readergirl12):

How many solutions does the equation have? 6-a=1/3(9a-18) A.Infinite solutions B.One solution C.No solution

OpenStudy (anonymous):

Okay. \[6 - a = \frac{ 1 }{ 3 }(9a - 18)\] \[6 - a = 3a - 6\] \[4a = 12\] \[a = 3\] \[Therefore,\] the answer is B... Because there is only one solution!

OpenStudy (anonymous):

Hope this helped! Have a great day! Also a medal would be much appreciated! Just click best response next to my answer. Thank You! @readergirl12